OpenStudy (anonymous):

Solve: -1/2x + 10 = −x + 12

OpenStudy (anonymous):

You have to isolate the constant and x to separate sides of the equals sign. To do this you should add x to both sides, and subtract 10 from both sides. Then either multiply or divide the number on x to make it 1, and do the same to the other side.

OpenStudy (anonymous):

-1/2x+10=-x+12 -1/2x+x=12-10 1/2x=2 x=4

OpenStudy (anonymous):

Thanks!!!